find the slope of the line 2x-5y=-16
Answer by checkley77(12844)

find the slope of the line 2x-5y=-16
-5y=-2x=-16
y=-2x/-5-16/-5
y=2x/5+16/5
Y=mX+b [m=slope=2/5] ans.